OpenStudy (anonymous):

eight less than a number doubled twenty more than the number what the number

OpenStudy (anonymous):

(Eight less than a number doubled) is (twenty more than the number). 2x - 8 = x + 20 x = 28 A

OpenStudy (anonymous):

@cedrick2654